### v1

# AI paintings vs. Human Paintings? Deciphering Public Interactions and Perceptions towards AI-Generated Paintings on TikTok

Abstract
With the development of generative AI technology, a vast array of AI-generated paintings (AIGP) have gone viral on social media like TikTok. However, some negative news about AIGP has also emerged. For example, in 2022, numerous painters worldwide organized a large-scale anti-AI movement because of the infringement in generative AI model training. This event reflected a social issue that, with the development and application of generative AI, public feedback and feelings towards it may have been overlooked. Therefore, to investigate public interactions and perceptions towards AIGP on social media, we analyzed user engagement level and comment sentiment scores of AIGP using human painting videos as a baseline. In analyzing user engagement, we also considered the possible moderating effect of the aesthetic quality of Paintings. Utilizing topic modeling, we identified seven reasons, including looks too real, looks too scary, ambivalence, etc., leading to negative public perceptions of AIGP. Our work may provide instructive suggestions for future generative AI technology development and avoid potential crises in human-AI collaboration.
Introduction
With the rise of various Generative AI technologies such as “Midjourney”, “stable diffusion” and “ChatGPT” in 2022, a vast array of AI-generated content (AIGC), such as images, texts, sounds, and even videos, has widely spread across major social media platforms. AIGC’s powerful functions and ease of use have gained the favor of many industries, particularly the art industry (Hitsuwari et al. 2023). Many famous artists, such as Sougwen Chung, Helena Sarin, Refik Anadol, etc., created advanced and groundbreaking artworks in cooperation with generative AI. As Sougwen Chung said [^1] https://zhuanlan.zhihu.com/p/149217204, “ As an artist working with these AI-generated tools, the prospect of artificial intelligence offers me a fresh perspective.” The AIGC has given the art industry a lot of new opportunities and provided artists with motivation.
However, the public has shown resistance towards genAI art. In December 2002, many painters uploaded images stating “NO TO AI-GENERATED IMAGES” [^2] https://arstechnica.com/information-technology/2022/12/artstation-artists-stage-mass-protest-against-ai-generated-artwork/ and called for a global boycott of AI-generated paintings (AIGP). This incident started because some artists felt that AI paintings had plagiarized their paintings without consent and their copyrights had been violated. This struggle lasted for nearly a year, and until now many painters are still calling for a ban on uploading AI-generated paintings on social media.
As AIGC circulates widely on social media, scholars have investigated how factors like text quality (Zhang and Gosline 2023) and AI usage (Messer 2024) affect viewers’ perceptions. However, many overlook AI’s painting capabilities and the aesthetic quality of AIGP on viewer participation and attitudes. Ragot et al. found significant cognitive bias in public perceptions of AIGP but did not consider Image Aesthetic Quality (IAQ) as a variable (Ragot, Martin, and Cojean 2020). Other studies (Bhandari, Chang, and Neben 2019; Shi, Huo, and Hou 2021) have shown that aesthetic quality significantly influences decision-making and perceptions. Therefore, we included IAQ as a potential moderating variable affecting user engagement.
This paper presented a comprehensive analysis of public perceptions towards AIGP based on user engagement, comment sentiment scores, and topic modeling. User engagement is a multidimensional concept that encompasses not just behavioral aspects (actions) but also emotional aspects (feelings) (Khan 2017; Hollebeek 2011). However, high levels of engagement can only show users’ interaction intention while cannot indicate completely positive perceptions. Considering this, we also adopted comment sentiment analysis to compare the emotional scores of comments between AIGP and human paintings for further perception analysis. Based on the above research motivations and feasibility, we formulated our research questions as follows:
RQ1: How are people’s intentions to engage with AIGP on social media platforms?
RQ2: How does the image aesthetic quality (IAQ) moderate people’s engagement intentions with AIGP?
RQ3: How are public sentiments toward AIGP? What factors motivate people’s negative attitudes toward AIGP?
To answer these questions, we used human paintings as the baseline and compared the differences in user engagement levels and comment sentiments between AIGP and human paintings. To conduct the comparisons, we compiled a dataset with 207 AIGP and 210 human-painting image-text videos and over 80,000 comments on these videos from TikTok. Image-text video is a format similar to PPT, containing only images and text, where users swipe left or right to switch between images in the video. To answer RQ1, We used a formula to evaluate the user engagement level for every painting video. To answer RQ2, we adopted improved-aesthetic-predictor (LAION-Aesthetics V2333https://laion.ai/blog/laion-aesthetics/) as the image aesthetic quality assessment (IAQA) model to assess the IAQ of each painting for further comparison. To answer RQ3, we leveraged Natural Language Processing (NLP) techniques for sentiment analysis and topic modeling. The findings revealed that people still express stronger interaction intentions when browsing human paintings and that the IAQ does not significantly affect engagement intentions. We also found that people have more negative attitudes toward AIGP than human painting. The results of topic modeling listed seven main reasons for the negative sentiment. Early analysis of the public perception of AIGP will help the ICWSM research communities understand needs and articulate designs and interactive technologies that reduce plausible societal harm.
Figure 1: User interfaces of TikTok
Background and related work
In this session, we briefly discuss the current background of AIGP technology development, what people think about the appearance of AIGP, and what the image aesthetic quality mentioned in this article is. We also provided loads of previous studies to support our research’s objective and necessity.
Social Attitudes toward Generative AI (Gen-AI)
With the rapid development of AI, it has become increasingly important to study societal attitudes towards Gen-AI. Many researchers have also studied the attitudes of various groups of people towards generative AI technology in recent years. Grassini S et al. (Grassini 2023) developed and validated an Artificial Intelligence Attitude Scale (AIAS) designed to assess public perceptions of AI technology. Vasiljeva T et al. (Vasiljeva, Kreituss, and Lulle 2021)found that attitudes towards AI varied widely across industries. A significant difference in attitudes towards AI was found between employees of organizations implementing AI solutions and those without the intention of implementing AI solutions. Pinto et al.’s (Pinto dos Santos et al. 2019) survey of 263 medical undergraduates found that undergraduate medical students were not concerned about AI replacing human radiologists, and they understood the potential applications and impact of AI on radiology and medicine. YILDIZ T et al. (YILDIZ 2023)found that language learners were delighted and favored using AI in the learning process. All in all, people seem to have different attitudes towards AI technology based on different usage scenarios and different populations.
AIGP, as the hot AI field of the moment, is developing at an astonishing rate, and the Paintings AI produces are already comparable to or surpass those created by humans. The technology not only brings us irreplaceable productivity and convenience (Xu 2019), but also the potential threat of indistinguishable deepfake problems (Bregler, Covell, and Slaney 2023), and Substitution of work  (Bendel 2023). Therefore, to determine the direction of future AI technological development and people’s needs, it is crucial and necessary to research the acceptance of artworks created by AI (Colton, Pease, and Saunders 2018). Many scholars have conducted social surveys on people’s attitudes towards AI-related work. Simon Colton (Colton 2008) was concerned about the perceived rejection of creative computers, but the evidence was unclear. Moffat and Kelly (Moffat and Kelly 2006) proved that significant bias for computer-generated music existed. Martin Ragot et al. (Ragot, Martin, and Cojean 2020) invited 565 participants to rate paintings created by generative AI tools and humans, respectively, and the results presented that the scores of human paintings were significantly higher than AI paintings. However, the Painting quality question in this study has not been solved, and the generative tools were incomplete. So, few studies constructed a complete process to analyze the attitude toward AIGP, primarily focusing on the painting artwork circulating on social media. In this research, we studied the social perception toward AIGP videos posted on social media and analyzed viewers’ sentiments from the interaction behaviors.
Image Aesthetic Quality Assessment (IAQA)
With the surge in visual artwork created by computers, we have witnessed AI-generated Paintings of varying quality (Ramesh et al. 2021). Meanwhile, it is increasingly challenging to filter and distinguish the chaotic and topic-unrelated images for some social media platforms. Therefore, how to efficiently and automatically evaluate the aesthetic quality of images spread on social media platforms has become an increasingly crucial problem.
Aesthetics, a significant branch of the visual arts, delves into exploring aesthetic categories such as beauty and its counterpart, ugliness (Zhu 2010). The aesthetic appeal of an image is evaluated based on universally accepted principles of photography. This appeal can be influenced by various elements, such as the strategic use of colors (Freeman 2007), the careful manipulation of contrast (Itten 1975), and the thoughtful arrangement of the image composition. Chatterjee and Leder et al. (Leder et al. 2004) introduced a five-stage, multi-level information processing model for image aesthetics. This model encompassed the perception of low-level information such as color, contrast, and complexity, the implicit integration of personal experiences and memories, explicit classification, cognitive mastery, and evaluation, culminating in aesthetic judgment and the generation of aesthetic emotions. However, computationally modeling this sequence for visual art images presents a significant challenge. Subsequently, Microsoft Asia Research Institute and Tsinghua University collaborated on a method to automatically differentiate photographs captured by professional photographers from those taken by customers (Tong et al. 2005). This study is widely recognized as the earliest research in the field of IAQA. The researchers employed a set of 21-class, 846-dimensional low-level global features to train a classification model for aesthetically categorizing the test images. Until 2014, with the emergence of AVA, a large-scale image aesthetic analysis dataset, image aesthetic analysis has been given an enormous boost based on deep learning technology (Jin et al. 2019). Lu X et al. (Lu et al. 2014) employed a dual-channel deep neural network to process local and global image blocks separately. Kong S et al. (Kong et al. 2016) developed the AADB aesthetic database, rated by human evaluators, and utilized deep CNN for aesthetic ranking and categorization of images. Chang KY et al. (Chan et al. 2019) created an aesthetic dataset, PCCD, and introduced a CNN and LSTM-based deep neural network capable of generating aesthetic linguistic comments on images. Jin X et al. (Jin et al. 2019) designed a novel deep convolutional neural network, ILGNet, for distinguishing between images of high and low aesthetic quality. Some famous datasets such as AVA (Aesthetic Visual Analysis) (Murray, Marchesotti, and Perronnin 2012), AADB (Aesthetics and Attributes Database ) (Kong et al. 2016) and CUHK-PQ (CUHK-Photo-Quality) (Ke, Tang, and Jing 2006) have collected billions of image samples. Therefore, using the aesthetic quality assessment models trained based on these datasets to evaluate the IAQ has been a good choice.
Data and Methodology
We integrated several quantitative methods to solve and answer the proposed research questions. We compared user engagement and comment sentiment between 207 short videos related to AIGPs and 210 short videos showing human paintings. Additionally, to identify prominent reasons for negative comments, topic modeling was adopted to analyze the 42199 comments for AIGPs and 39763 comments for human paintings, respectively.
Hypotheses
Previous research has indicated that people generally hold negative attitudes towards AI-generated content, particularly in the realm of artworks. Therefore, our Hypotheses 1 and 3 are also based on this perspective, but we have narrowed the scope to AI-generated paintings (Wu et al. 2020). Meanwhile, visual aesthetic quality has been shown to be a significant factor in promoting user behaviors such as downloading and rating items on the Internet (Bhandari, Chang, and Neben 2019). Consequently, based on prior studies, we also posit that the aesthetic quality of AIGP positively impacts user engagement willingness (H2). To answer the research questions, we proposed the following three hypotheses:
H1: People are more willing to interact with human paintings than with AIGP on social media platforms.
H2: AIGPs with high aesthetic quality scores lead to higher user engagement.
H3: People express more negative emotions in comments on AIGP than human paintings on social media.
Research Site
TikTok is the fastest growing social network in the post-pandemic era, allowing users to publish and share short videos ranging from 15 to 60 seconds. It had 1506 million downloads in 2021 and has been the most downloaded social media application globally, surpassing Instagram, which had 1048 million downloads (Barta et al. 2023). In 2022, the average daily time spent on the TikTok mobile app by users worldwide was almost equivalent to that of Facebook and Instagram combined (Zhang et al. 2023). There are three main reasons for choosing TikTok as the survey platform for this study. First, as of December 15, 2023, videos under the tags related to AI painting on TikTok, such as “aiart”, “midjourny”, and “stable diffusion”, have received 10.7B, 7.5B, and 2.3B views. This not only indicates that a large amount of content related to AIGP has been widely disseminated on the TikTok platform, but also shows that users have a high degree of motivation and willingness to participate in discussions about AI-related topics. Second, based on TikTok’s recommendation algorithms, the platform analyzes users’ viewing preferences and directly pushes videos to them on the “For you” page (Zhao 2020). In this process, the viewer has no idea what the next video is about, thus avoiding the influence of users’ utilitarian participation on the video views in the experimental data. According to the Authentic Self Theory, individuals are more likely to exhibit genuine reactions when confronted with random events because they have limited time to prepare or control their responses (Vannini and Franzese 2008). Therefore, when a video with uncertain content like the one in TikTok suddenly “pops up” in front of the user, the user’s reactions are more likely to be genuine. Last but not least, according to TikTok’s platform management guidelines, all videos involving AI-generated content must be labeled, and videos without labels will not pass the platform’s review. Therefore, compared to other social media platforms such as Reddit, Instagram, and Twitter, TikTok is undoubtedly the most appropriate choice.
Data Collection and Preprocessing
Figure 2: Video Filtering Flowchart
In our preliminary survey, we discovered that paintings on TikTok are presented in two formats: short videos and image-text videos. In short videos, many creators share their motivation for creating, record the creative process, or provide step-by-step painting tutorials. In these videos, user engagement is influenced by many factors unrelated to the artwork itself, such as the creator’s voice and behavior, the video’s length, etc. (Yang et al. 2022), which means that it is difficult to capture people’s actual attitudes towards the two types of paintings from these videos. Therefore, we selected image-text videos as our experimental objectives, which are similar to PPT format and contain only images and text. Users swipe left or right to switch between images in the video. In this kind of video, creators act more as sharers than narrators, only sharing their creations or AI-generated paintings. Thus, user engagement depends entirely on their understanding and feelings towards the artwork, unaffected by the creator’s value output. Therefore, choosing image-text videos for comparative analysis perfectly aligns with our research theme, making our experimental results more scientific and reliable.
Fig.2 a shows the initial steps for video collection and preprocessing. We created a new account on the TikTok video-sharing social media platform specifically for this study. By searching for tags related to AIGP and human paintings on TikTok, such as “Stable diffusion”, “Midjourney”, “AI paintings”, “Procreate” and “SAI” and scraping the web, we selected 2667 AI-labeled image-text videos and 1729 human painting image-text videos as of November 15, 2023. Considering videos with low views might not be representative, and some creators might restrict comment permissions, we excluded videos with fewer than 10k views, irrelevant content to the painting, and less than five comments. This refined our dataset to 778 AIGP videos and 935 human painting videos. To avoid favors or bias from users toward specific creators and repetitive Paintings, we adopted the same-author avoidance principle, ensuring no more than five works from the same author in our sample and excluding videos with repetitive Paintings within the Painting. Finally, after filtering based on these basic rules, we obtained 207 AIGP image-text videos and 210 human painting image-text videos.
Figure 3: Image Aesthetic Quality Schematic
User Engagement Analysis (RQ1)
In this work, we define user engagement as the interactivity between users and videos (O’Brien and Toms 2008), reflecting the audience’s genuine interest and interaction with the videos. This is considered an indicator of the level of trust users have in a creator and their skill level. We adopted a formula for calculating TikTok user engagement on the HypeAuditor platform, an AI-driven analytical and discovery tool for Instagram, YouTube, TikTok, Twitter, and Twitch. It helps to find relevant influencers, understand their audience, increase the ROI of advertisers, and maintain the authenticity of influencer marketing. Many previous studies on social media data analysis have used this tool(Rogers ; Gil-Muñana and Sáez-Linero 2023). The formula for TikTok user engagement on Hype is $[(numberoflikes+numberofcomments+numberofshares)/numberofviews]X100\%$  [^4] https://hypeauditor.com/blog/what-is-tiktok-engagement-rate-why-brands-should-take-note/.
Image Aesthetic Quality Assessment (IAQA) (RQ2)
Video quality is an important control variable in this paper, and its impact on user engagement has been confirmed by many previous studies(Diallo, Fieau, and Hennequin 2014). So, in the process of sample preprocessing, we also considered the video quality and the flowchart is presented in Fig. 2b. Since image-text videos only contain Paintings and text, all videos have the same page layout and presentation method, so here, video quality is equivalent to the aesthetic quality of the Paintings in the videos. Firstly, we downloaded all the Paintings from the two types of videos, removed duplicate Paintings within the same video, and obtained 2383 AI paintings and 1153 human paintings. In our research, we used the improved-aesthetic-predictor (LAION-Aesthetics V2) [^5] https://laion.ai/blog/laion-aesthetics/ as the IAQA model. The model’s training process is divided into two steps: embedding the input images using Clip ViT/14 (Radford et al. 2021) and then using MLP to train the image data from three well-known datasets, namely, 176,000 image-text pairs from SAC (Pressman, Crowson, and Contributors 2023) dataset, 15,000 image-text pairs LAION-Logos dataset with aesthetic scores from 1 to 10, and 250,000 photos from the AVA dataset with aesthetic scores from 1 to 10. The data after aesthetic quality scoring of 2.73 billion images in the LAION 5B dataset (Schuhmann et al. 2022) by this Predictor trained Stable Diffusion V1666https://github.com/CompVis/stable-diffusion/tree/ce05de28194041e030ccfc70c635fe3707cdfc30#stable-diffusion-v1 , one of the most professional AIGC models currently. Therefore, this model’s comprehensive reliability and accuracy in assessing image aesthetic quality are likely the best now available.
To control the quality of the two types of videos at the same level and avoid the interference of extreme values, we excluded videos with less than $4.5andmorethan$6.5 average image aesthetic quality scores, ultimately obtaining 207 AIGP videos and 210 human painting videos. The independent sample T-test results (Table 3 in Appendices) also confirmed that there was no significant difference in IAQA values between the two types of videos (t=1.499, P = $0.135>$0.05). We then divided the videos into four quality intervals: Q1: (4.5,5.0], Q2: (5.0,5.5], Q3: (5.5,6.0] and Q4: (6.0, 6.5]. We illustrate some Paintings from different aesthetic intervals in Fig.3. By counting the total number of views of the two types of videos, we find that their total number of views has reached hundreds of millions, which indicates that a large number of users are viewing similar videos on TikTok. Therefore, our sample data is universal and reliable.
Sentiment Analysis and Topic Modeling (RQ3)
In this study, sentiment analysis was conducted on comments from two types of videos to reflect the genuine inner feelings of users, thereby revealing people’s real attitudes towards AIGP on social media. Fig.5 shows the flowchart of comment sentiment analysis and comparison. We extracted all comments from the sample videos, resulting in 392,812 comments for AIGP videos and 63,445 for Human painting videos. First, we used the open-source model ’Hello-SimpleAI/chatgpt-detector-roberta’ (Guo et al. 2023) on Hugging Face to detect if any comments were generated by bots, and then we removed those comments. Comments that were only “@” or had received no likes were removed too. After filtering, 42,199 comments from AIGP videos and 29,763 comments from Human painting videos were retained. These comments were then inputted into a sentiment analysis model named “distilbert-base-multilingual-cased-sentiments-student”  [^7] https://huggingface.co/lxyuan/distilbert-base-multilingual-cased-sentiments-student. It’s an open-source model based on distilbert (Sanh et al. 2019) training designed for multilingual applications, capable of recognizing twelve languages and interpreting sentiments as positive, neutral, or negative. As of December 30, 2023, it has been downloaded over 7 million times on Hugging Face. Since the model recognizes only twelve languages, ChatGPT 4.0 was used to translate the comments outside of the twelve languages into English. The model’s output categorized positive sentiment comments as 1, neutral comments as 0, and negative comments as -1. Since the final results did not conform to a normal distribution, the Wilcoxon test was chosen to compare the sentiment scores of the two types of videos.
For topic modeling, we initially filtered the comments on AIGP videos to extract all negative comments, resulting in 19,071 comments. These comments were subjected to topic modeling in this study. After comparing several mainstream topic modeling methods, such as Latent Dirichlet Allocation (LDA), Latent Semantic Analysis (LSA), and Non-Negative Matrix Factorization (NMF), we opted for the NMF model because our data contained amount of short texts but not long documents (Gamage et al. 2022). We trained the NMF model for different numbers of topics from k = 1 to 15, and we calculated the mean topic coherence across all topics to determine the optimal number of topics automatically (Figure. 6 in Appendices). Semantic coherence involves examining the co-occurrence of words within a document to ensure that the topics generated by the model possess semantic cohesion  (Wang and Lu 2023), a mainstream method for assessing topic quality. When k = 1, semantic coherence was the highest, but a single topic could not accurately summarize our collected comment data. Hence, we chose 14 topics for the final training phase, as this number demonstrated a relatively high level of semantic coherence.
Results
Table 1: The user engagement calculated by traditional formula. The normality (N) of the data was assessed using the Shapiro-Wilk test. We used an independent T-test if the difference is normally distributed and Wilcoxon test otherwise.
Measures Normality Test types Items Mean Std W p
User engagement N Wilcoxon test AIGP videos (N = 207) 0.12 0.07 -13.774 0.00***
HP videos (N = 210) 0.23 0.08
Table 2: A comparison of the comment sentiment scores. The normality (N) of the data was assessed using the Shapiro-Wilk test. We used an independent T-test if the difference is normally distributed and the Wilcoxon test otherwise.
Measures Normality Test types Items Mean Std W p
Sentiment scores N Wilcoxon test AIGP videos (N = 207) 0.3 0.93 -36.725 0.00***
HP videos (N = 210) 0.53 0.83
Informed by our hypotheses, all tests in the study were conducted using one-tailed comparisons. The analysis of our data and manipulation checks utilized independent T-tests, Wilcoxon tests, or one-way ANOVA, depending on whether the data distribution was normal. Our normality tests on user engagement for both types of videos revealed that the data did not conform to a normal distribution. As a result, we employed the Wilcoxon test to compare the user engagement of the two sample types, as detailed in Table 1  (Hu et al. 2022). Table 2 compares the comment sentiment of two types of videos. The results supported hypotheses H1 and H3, indicating that people’s user engagement (W = -13.774, P = 0.00) and sentiment scores for comments (W = -36.725, P = 0.00) on human painting videos were significantly higher than those for AIGP videos. These findings address RQ1 and RQ3-part1 posed in the paper.
Regarding the impact of video aesthetic quality on user engagement, the videos’ average aesthetic quality was divided into high, medium, and low ranges according to the LAION 5B  (Schuhmann et al. 2022) divisions for IAQA: (5.5, 6.5], (5.0, 5.5], and (4.5, 5.0]. We then assessed the impact of these quality ranges on user engagement, as detailed in Table 4 in Appendices. As this involved comparisons among three samples, we conducted one-way ANOVA tests on user engagement corresponding to different quality intervals for AIGP videos and human drawing videos, respectively. The results indicated that image aesthetic quality does not significantly affect user engagement (F = 2.056, P = 0.131; F = 1.663, P = 0.192). Moreover, we found that the difference in user engagement between the two types of videos in different quality intervals is also very close to each other. The difference in high IAQ is 0.123, in middle IAQ is 0.122, and in low IAQ is 0.133. Hence, the experimental results indicate that hypothesis H2 is supported, and RQ2 posed in the paper is answered.
RQ1: How are people’s intentions to engage with AIGP on social media platforms?
As seen in Table 1, when the quality of the video content is the same, people are more willing to interact with human painting videos than with AIGP videos. Because our experimental data collection and pre-processing process has eliminated the interference of most other variables, in such a situation, the response made by the user to the video can almost be equated with the user’s real attitude towards the paintings. So, the users’ behaviors and decisions are solely based on the paintings, such as liking, commenting, favoriting, and sharing. We can see that although there is a lot of AIGP-related content on social media and there are many people who have viewed AIGP works, users still prefer to browse and interact with human paintings on social media platforms.
Stronger Desire to Share AIGP
We calculated the like, comment, favorite, and share ratios of each video based on views in Table 5. We found that the like, comment, and favorite ratios for human painting videos significantly exceed those for AIGP videos, particularly in the comment ratios of 0.001 (AIGP) and 0.003 (HP). These data corroborate our previous results on user engagement comparison, and also show that using machine learning methods to calculate user engagement is reliable. Interestingly, the share ratio of AIGP videos is much higher than that of human painting videos. Based on our speculation, the possible reason is that the ability of AI painting breaks people’s perception of AI technology, and the speed of its development is also shocking to people who want to share it and discuss it with their friends. However, as with comments, we are not clear about users’ motivations for sharing, which can be positive signals or negative. Therefore, we analyzed the comment motivation specifically in a later section.
RQ2: How does the image aesthetic quality (IAQ) moderate people’s engagement intentions with AIGP?
We compared the impact of different aesthetic quality intervals of Paintings on user engagement for both video types, and the final results found no significant correlation between the two. This suggests that when browsing different artworks on social media, people may be more concerned with the intrinsic meaning and value of the paintings than their glossy appearance. Moreover, we noticed that the engagement with human painting videos increases slightly as aesthetic quality decreases. Observing the comments on human painting videos, there appears to be a lot of positive social support for the creators, such as when amateur painters upload less-than-perfect paintings and drawings on TikTok. Interestingly, many users offer encouragement rather than criticism. Some examples are as follows:
The Drawing Looks Better Than The Original!
your art just gets better and better I’m amazed!
YOUR OLD ART IS SO PRETTY BUT YOUR CURRENT ART IS EVEN BETTER!! You’ve definitely improved I’m obsessed with your style.
Such emotional support (e.g., encouragement, intimacy, compliment, and empathy) can give individuals a sense of respect and achievement, strengthening community ties and member engagement (Wang, Kraut, and Levine 2012). At the same time, viewer behaviors may also be influenced by the herd effect (Yen et al. 2023), following the content of the majority’s comments and involuntarily giving a steady stream of support to the art creator. This emotional value then stimulates the creator’s willingness to create. So this may be one of the reasons why the lower the aesthetic quality of the Painting, the higher the engagement of human painting videos. But we don’t find such a large-scale phenomenon in AI painting videos.
RQ3: How are public sentiments toward AIGP? What motivates people’s negative attitudes toward AIGP?
After comparing the comment sentiment scores of both video types, we found in Table 2 that human painting videos have significantly more positive comments than AIGP videos. It is undeniable that the number of positive comments about both types of paintings is higher than the number of negative comments, and this data suggests, to a certain extent, that many people have already accepted that AI painting exists. Some AI-created artworks are also well-liked. However, compared to human artworks, some people tend to harbor more negative emotions when discussing AIGP content on social media. Overall, people still prefer to see human artworks on social media platforms rather than AI-generated ones.
This section addresses RQ3-part 2 by identifying seven types of negative comments and reasons for reluctance to interact with or express negative sentiments toward AIGPs on social media. The topic modeling revealed 14 common topics within the dataset, among which the top 7 topics with the highest proportion are listed in Table 6. To avoid the influence of positive comments, all topics were extracted from 19,071 negative comments about AIGP posted on TikTok.
It is too real/appalling to be unacceptable
After analyzing the topics, we found a slight overlap between Topic 1 and Topic 7, and we discussed these two topics together. These topics reflected that the reason for people’s negative attitude toward AIGP is that AIGP has become too realistic and is perceived likely to affect our lives in the future. Indeed, the advancement of AI technology has become increasingly unpredictable, and we are facing a dilemma that it is difficult to distinguish between human paintings and AIGPs. This dilemma will become increasingly challenging to break free from, along with the development of AI technology. One controversial incident involved using a deepfake voice in the documentary film “Roadrunner” about celebrity chef Anthony Bourdain (Gamage et al. 2022). The film incorporated an artificially artificial voice to simulate Bourdain speaking, which he never originated or consented to. This means that we may not even know when others have privately used our voices and likenesses. So when users are browsing these AIGPs, it’s hard for them not to conjure up Paintings of the day when we can no longer tell which ones are AI-generated and which ones are human’s own. Examples are given below:
If you don’t say AI, you won’t know AI.
I am afraid that the unique feeling of AI will be eliminated by refining some of the real ones.
Finally, it’s getting harder for AI to discriminate…
They all express deep concern about the development of AI and the future of deepfake. This is, therefore, one of the most important reasons for users to oppose AI, accounting for 12.82 percent.
I feel like I’m in a spiral of self-contradictions
Topic 2 expresses the idea that when many users see AIGPs, they are indeed attracted by the fine character portrayal, rich color palette, and other artistic features, but when they see the AI tags underneath the video, they start to hesitate and sway from side to side or even feel disgusted and ashamed of the paintings, like these typical comments:
This looks so gorgeous! Shame it’s AI Art tho
Disappointed when I realised it’s AI art, stop taking from the community and faking art
the way my smile went away when i noticed its AI
Users have already recognized the painting and think it is really good artwork, but considering the AI technology behind the artwork, they turn from a positive to a negative attitude. Therefore, the keywords “cute” and “interesting” in this topic do not represent a positive emotion but mockery and self-congratulation.
I am not acceptable for AI paintings, and it’s not art.
Topic 3 illustrates that some users have a cognitive bias towards AI art, a tendency or preference based solely on subjective perceptions and thought patterns and that other factors do not influence this bias. This cognitive bias has also been confirmed in previous research on AI paintings (Ragot, Martin, and Cojean 2020). Such a phenomenon is similarly confirmed in this paper. In their perception, AIGP can never be called art, as this comment says:
Sadly AI art isn’t art, this is just slide after slide of dogshit
if it’s made by AI it’s not art
Thus, the negative attitude of this category towards AIGP is entirely unjustified, at least on social media, where they do not indicate the reason for their distaste for AIGP. Furthermore, this unjustified non-acceptance may not be limited to reviews on AIGP but on all things related to AI, such as:
I hate ai soo much
The fact that its AI generated makes it boring and empty
It is not good enough to meet my expectations.
Many AIGPs are generated based on real people, anime characters, or classic scenes. The concept of Topic 4 is that many people think that AI-generated characters in Paintings don’t live up to their expectations or don’t capture some of the character’s traits, which leads to their negative attitudes about AIGP. Some comments are given below:
i think AI did kyo n yuki dirty nooo
AI inconsistency when it comes to skin colour
These comments highlight the conflict and mismatch between the AI-generated characters and the Painting of the character in people’s minds. Some avid fans of anime characters consider that AI ruined their favorite characters, and as a result, they show strong resentment and anger in the comments.
It is so scary and weird.
There are two meanings we can get from Topic 5. Firstly, the word “dark” indicates that there may be more dark and eerie types of AIGPs on TikTok, and this style of drawing makes people feel uncomfortable. On the other hand, AI-generated technology has always had some complex technical difficulties, such as the drawing of human limbs and skin color:
AI is pretty inconsistent when it comes to darker skin colors and the hair accuracy… shame but it has nothing to do with me
Some of the keywords in this topic, like “potter” and “hand” indicate that some human-themed AIGPs on social media have weird limbs or postures. These Paintings are likely to horrify and disturb users, leading to negative comments.
It’s plagiarism
The reason for the negative attitude towards AIGP, as explained in Topic 6, is that AIGP’s creative process is perceived as a form of plagiarism, and therefore, people feel cheated. Some typical examples as follows:
idk if im impressed or scared like these ai art are STEALIN artists art omd
many AIs that seek to draw the human, what they do is plagiarise the art of a list of authors, is just coming out.
AI art is theft
People think that why AI can generate high-quality paintings because of plagiarising the artists’ paintings, which are not thought out by themselves, and therefore, AI painting should not be accepted and spread.
Discussion
Humans and AI have become inextricably linked. Many examples and studies have shown that human-AI collaboration will lead to better creativity in many industries (Maerten and Soydaner 2023). However, according to some negative news, public feelings seem to be ignored, while generative AI is increasingly widely used for creative activities. Such discussions about public consciousness and awareness when living with AIGC are very critical, which ensure the competency of AI implemented responsibly and judiciously (Sengupta, Srivastava, and Mcneese 2024). Therefore, the purpose of this study is to analyze users’ discussion content and motivations toward AIGP on social media.
Hesitation and Tentativeness in Interaction with AIGP
According to Nadarzynski’s study of the public response to AI-led chatbot services, “AI hesitancy” was prevalent during users’ interaction with chatbots  (Nadarzynski et al. 2019). It demonstrated that a substantial number of people may feel hesitant to use AI modules, considering the accuracy and security of these services. By comparing the degree of users’ interaction between the AIGP video and the human painting video, we found that people still feel dithery when responding to the AIGP on social media, although the interaction does not involve services or device usage. This low level of willingness to engage is reflected in the likes, favorites, and comments on AIGP-type videos. Taking into account possible existing technical reasons for generative AI, we also controlled the moderating variable of IAQ. However, the moderating role of IAQ is not significant. This finding reinforces the fact that there may still be an unacceptable attitude and mistrust in regard to the generative AI technology itself. Introducing the Fear theory into the AI-related research, several studies have acknowledged that new technology is frequently perceived as a potential threat due to the risks it may pose to individuals or society as a whole. (Osiceanu 2015; Khasawneh 2018). Regarding AI specifically, we are currently undergoing a shift towards cities and societies that are governed by various artificial intelligences (Cugurullo and Acheampong 2023). However, the overall risks of this transition are still not fully understood (Yigitcanlar et al. 2020). Based on these previous unresolved questions, this paper contributes interesting empirical findings to research related to generative AI risks and social attitudes. We derived seven topics from users’ comments on AIGP to uncover why negative attitudes still exist. Among the topics, a negative attitude due to the fear of AIGP was confirmed in this study. We also find that ambivalence and looking too real were typical reasons that led to negative attitudes.
Influence of AIGP to Art
There is no doubt that AIGP has had a significant impact on the art industry, as can be seen from the results of topic modeling, with Topic 1 and Topic 7 all highlighting the fact that today’s AIGPs are so similar to human paintings that we cannot distinguish some of the high-quality AIGPs from human paintings. On December 6, 2023, a post made it to the top of Weibo (one of China’s most famous social media platforms, similar to Twitter) hotlist, which was along the lines of the mascot for the Spring Festival Gala announced by China Central Broadcasting Television (CCTV) being thought to have been generated by AI. Some users even posted long posts analyzing various details to prove that it was generated by AI. Although an official announcement was made to reply to the challenge and the designer’s original design was released, users still firmly believed it was AI-generated and accused of this behavior. As a result, the emergence of AIGP has added many burdens to artists. As far as painters are concerned, they not only have to design creative and high-quality artworks but also need to worry whether their works will be labeled as AI and blamed. Especially, some high-level painters have absolute confidence that their paintings can be better than AIGPs. Still, they are afraid that the paintings they spent a lot of time and energy to create will be regarded as AI-generated. Therefore, this may be one of the reasons why AIGP is likely to cause many excellent artists to lose their creative drive and support anti-AI events.
Implications for social media platform design
This paper explored the effect of image aesthetic quality on user engagement, and the experimental results confirmed that the image aesthetic quality of paintings on TikTok does not have a significant impact on users’ interactions. We found that users showed more robust emotional support and willingness to interact with some human paintings with relatively low aesthetic quality. This indicates that when browsing paintings on social media, users prefer to see the creator’s continuous improvement process, even if the work is not perfect. However, in AI-generated paintings, we rarely see such emotional support. Therefore, this finding can provide some valuable insights into the study of social support in online communities, an important topic in social media research.
In addition, we analyzed and compared the sentiment of comments on AIGP and human paintings. The results showed that positive comments generally dominated people’s comments on both types of paintings, but the rate of negative comments on AIGP videos was significantly higher than that on videos of human paintings. It means that there are still a lot of people who have a negative attitude towards AIGP, and many people are likely unwilling to see AI-generated Paintings on social media platforms. Therefore, in order to further explore what makes people hold negative attitudes towards AIGP content on TikTok, we conducted topic modeling of negative AIGP comments and finally identified seven main reasons. These reasons provide suggestions and guidance for optimizing future AI-generated models, from which we can analyze people’s preferences to determine which AIGP to put on social media to attract users’ attention and promote an intimate connection between the platform and users.
One of our findings is that people prefer viewing human-created paintings. This suggests that platforms can moderately reduce the promotion of AIGP and encourage artists to upload their own work, which can increase user engagement. On the other hand, we find that people’s comments on AIGP videos express more negative emotions, which suggests that platforms may need to pay more attention to comments related to such videos to avoid vicious incidents.
Limitations and Future Work
This study uses user-generated content (UGC) on the TikTok platform as the source of experimental data, so data collection was done on a single social media platform. However, whether there is a difference in the perception of AIGP among users of different platforms, such as Twitter, Facebook, Reddit, etc., is an area not considered in this paper. However, it is undeniable that the number of active users on TikTok is still the highest of all social media platforms, and all AIGP videos have to be tagged with “ai”, which very few platforms do. We also tried to spread our survey to other mainstream social media. Still, we found that many platforms are not open and transparent about their data, such as the number of video videos or post views, which is an essential measure of user engagement in this experiment. Thus, this can cause significant difficulties in collecting data for our experiment, and the accuracy of the data cannot be guaranteed. At the same time, we did not focus only on numerical data on the platform. We also processed and analyzed tens of thousands of comments, which reflect the actual attitudes and emotions of platform users towards the two types of paintings. Consequently, integrating these several methods and data can guarantee the rigor of the experimental results. To further validate our findings, our future work will no longer be limited to social media platforms, and we hope to investigate and analyze these issues in this paper in more detail through questionnaires or real-life interviews.
Conclusion
Because of TikTok’s huge user groups and development speed far exceeding that of other platforms such as Facebook and Twitter, people’s interactions and perceptions towards AIGP on the TikTok platform can reflect a social phenomenon to a certain extent. Therefore, this paper investigates the user engagement and emotional tendency of comments on AIGP on social media based on the TikTok platform and compares it with human drawings. Finally, it is concluded that people are more willing to interact with the content of human paintings on social media and have higher positivity towards their comments. We also considered the potential moderating role of the IAQ in human-AIGP interactions, and the results did not support this hypothesis. Finally, seven main themes derived from the topic modeling, including looks too real, ambivalence, etc., indicated why some people still have negative perceptions towards AIGP.
Appendix A Appendices
Measures Normality Test types Items Mean Std Video views t p
IAQA Y Independent T-test AI videos (N = 207) 5.48 0.46 259088700 1.499 0.135
HP videos (N = 210) 5.41 0.48 40870900
Table 3: A comparison of IAQA between AIGP videos and human painting (HP) videos. The normality (N) of the data was assessed using the Shapiro-Wilk test. We used an independent T-test if the difference is normally distributed and the Wilcoxon test otherwise.
Measures Test types IAQA AIGP videos (N = 207) Human painting videos (N = 210)
Sample
size
Mean std F P
Sample
size
Mean std F P Difference
User engagement (Formula)
One-way
ANOVA
(5.5,6.5] 109 0.13 0.068 1.788 0.17 84 0.23 0.072 0.016 0.984 0.1
One-way
ANOVA
(5.0,5.5] 66 0.11 0.069 80 0.22 0.081 0.11
(4.5,5.0] 32 0.12 0.069 46 0.22 0.09 0.1
Table 4: A comparison of the user engagement from different video quality intervals. We used the One-way ANOVA test for three samples.
Items
Like-to-view
ratio
Comment-to-view
ratio
Favorite-to-view
ratio
Share-to-view
ratio
AIG videos (N=207) 0.10107 0.00101 0.02057 0.00224
HP videos (N=210) 0.19388 0.0029 0.02779 0.00081
Table 5: Like ratio, comment ratio, favorite ratio, and share ratio
Topics Description Keywords Proportion
Too real (Topic 1) The AI-generated images look like authentic artwork created by true painters. like, looks, artist, sick, cute, real, bad, feel, movies, girl, style 12.82%
Ambivalent (Topic 2) Users shift their behavior from praise to criticism when they see it was generated by AI, although it is gorgeous. damn, hate, shame, fuck, cute, insane, great, disappoint, interesting, cool, pretty 12.72%
Unacceptable (Topic 3) The complete lack of recognition and acceptance of AI-generated art ai, art, generated, crazy, real, fuck, hate, sad, bad, shit, program 11.43%
Not meet expectations (Topic 4) The AI-generated image content did not meet the user’s expectations. damage, hand, emotional, hear, remix, eyes, expecting, brain, think, forget, remember 8.99%
Scary (Topic 5) The AI-generated images are horrifying and irrational. dark, scary, souls, minecraft, wars, hand, terraria, style, fortnite, potter, art 8.53%
Stealing (Topic 6) Users think the AI stole the artist’s work, so they feel cheated. did, use, make, theft, dirty, discord, program, wrong, cheating, steal, generate 7.65%
Appalling (Topic 7) AI painting ability is so strong that it has surpassed human painters. hard, crazy, pic, insanely, incredibly, unbelievably, side, hit, monkey, doom, shit 6.50%
Table 6: Results of topic modeling
Figure 4: lowchart for calculating and comparing user engagement
Figure 5: Sentiment Analysis Flowchart
Figure 6: Mean Coherence of Comments